### v1

# Discovering causal relationships between time series with spatial structure

Abstract
Causal discovery is the subfield of causal inference concerned with estimating the structure of cause-and-effect relationships in a system of interrelated variables, as opposed to quantifying the strength of causal effects. As interest in causal discovery builds in fields such as ecology, public health, and environmental sciences where data is regularly collected with spatial and temporal structures, approaches must evolve to manage autocorrelation and complex confounding. As it stands, the few proposed causal discovery algorithms for spatiotemporal data require summarizing across locations, ignore spatial autocorrelation, and/or scale poorly to high dimensions [29, 13]. Here, we introduce our developing framework that extends time-series causal discovery to systems with spatial structure, building upon work on causal discovery across contexts and methods for handling spatial confounding in causal effect estimation [10, 25]. We close by outlining remaining gaps in the literature and directions for future research.
Keywords: latent spatial confounder, conditional independence, causal discovery
1 Introduction
The identification of cause-and-effect relationships from observational data has always been at the center of scientific research [33]. Statistical methods with this explicit goal have only emerged in the last three decades, however, and are still catching up to the complexity of practically available data [6, 11]. Spatiotemporal data pose particular difficulty for causal discovery methods due to autocorrelation between observations and with latent confounders that may mask or distort relationships [29]. In this short communication, we posit that existing causal discovery algorithms for independent and identically distributed time series can be adapted to time series with spatial structure by leveraging the theoretical results of causal discovery across contexts and latent spatial confounding in causal effect estimation [44, 25, 10].
2 Causal discovery
The graphical causal approach proposes that systems can be represented by graphs in which causal variables are connected to their effects by directed edges (arrows) without forming feedback loops [59, 32]. Graph structures imply certain properties of a system’s joint probability distribution and inform the fitting of structural causal models [33]. Structural causal models (SCMs) are formed of a hierarchy of equations wherein each variable is a function of its direct causes, or parents, $\mathbf{pa}_{i}$ in the system’s graph as well as some variable-specific unmodelled error $u_{i}$:
$x_{i}=f(\mathbf{pa}_{i},u_{i})$
Graphs and associated SCMs can be used to evaluate counterfactual scenarios, make predictions, or find appropriate variables to control for when estimating a causal effect.
The graphs underlying SCMs are often defined by subject experts or as a representation of a specific hypothesis [38, 28]. In either case, some confirmation bias may be present as only hypotheses previously considered in the literature and/or by the modeller may be put forth. There are also instances where no good hypotheses exist a priori to describe a system and exploratory analysis must be undertaken, motivating estimation of a causal graph from observational data.
Score-based causal discovery methods estimate a best-fit graph by minimizing a causally-relevant loss function across the space of possible graphs [5, 18]. Constraint-based methods iterate through all pairs of variables, testing their independence conditional on other variables in the system, and estimate cause-and-effect relationships by eliminating structures that are inconsistent with conditional independencies and/or do not meet assumptions [56, 33]. We focus on constraint-based methods here as they do not impose any inherent parametric or functional assumptions [11, 6].
2.1 Constraint-based causal discovery
The PC (Peter-Clark) algorithm was proposed in 1991 as the first asymptotically correct constraint-based causal discovery algorithm that could handle more than a few independent and identically distributed variables [55]. In the decades since, the logic of the PC algorithm has been a popular jumping off point for extensions [41, 47, 13, 58, 49].
For PC and PC-derived algorithms, the goal is to estimate the placement and direction of edges in the causal graph of a system of variables. Given the $n\times p]$ system of variables $\mathbf{X}=\{X_{1},...,X_{p}\}$, the algorithm first supposes a fully connected graph with undirected edges between all pairs $X_{i},X_{j}\in\mathbf{X},i\neq j$. It then searches for separating sets $\mathbf{A}_{ij}\subseteq\mathbf{X}\setminus\{X_{i},X_{j}\}$ such that $X_{i}\mathrel{\perp\!\!\!\perp}X_{j}\mathrel{\,|\,}\mathbf{A}_{ij}$. Where (conditional) independence is found, or, equivalently, when such a set $\mathbf{A}_{ij}$ exists, the variables $X_{i},X_{j}$ must not be causally connected and the edge between them is deleted.
Once the placement of edges has been determined, they are directed according to the following constraints. Where dependence between otherwise (conditionally) independent variables $X_{i}$ and $X_{j}$ is introduced by conditioning on a mutually adjacent variable $X_{k}$, a collider or common cause is indicated and edges are directed $X_{i}\rightarrow X_{k}\leftarrow X_{j}$. Once colliders have been determined, remaining undirected edges are directed as possible to avoid inducing feedback loops/cycles and so as not to create false colliders. Edges may remain undirected when direction is not resolvable according to the constraints [33].
The PC algorithm makes the following assumptions [55, 33]:
  • Faithfulness (also known as the causal Markov condition): Conditional independence relationships faithfully correspond to causal structures
  • Sufficiency: All relevant variables are observed
  • Stability: Relationships are consistent across observations
Various extensions relax the faithfulness [41], sufficiency [58], and/or stability [13, 49] assumptions.
2.2 Conditional independence testing
In addition to its assumptions, constraint-based causal discovery hinges on consistent, accurate conditional independence testing. Algorithms tend to be independence test-agnostic, however, and rarely enforce the use of any particular test or make claims of asymptotic correctness beyond the “oracle” setting where independence tests make no errors [47, 36, 11, 6].
Some of this confusion is due to the nature of the problem. Shah & Peters proved in 2020 that there is no nontrivial test that can distinguish between the null and alternative hypotheses of conditional independence and dependence with consistent power greater than the significance level [52]. Restrictions must be placed on the null to ensure power at any alternative.
In the same paper, the authors propose the generalized covariance measure, or GCM, as a practical approach to testing conditional independence. The null hypothesis of this test is that $X_{i}$ and $X_{j}$ are independent given $\mathbf{A}$:
$\begin{split}\text{H}_{0}:\,&X_{i}\mathrel{\perp\!\!\!\perp}X_{j}\mathrel{\,|\,}\mathbf{A},\,\text{or, equivalently, }\\ &\,\mathbb{P}(X_{i},X_{j}\mathrel{\,|\,}\mathbf{A})=\mathbb{P}(X_{i}\mathrel{\,|\,}\mathbf{A})\mathbb{P}(X_{j}\mathrel{\,|\,}\mathbf{A})\end{split}$
The GCM test approximates these conditional distributions $\mathbb{P}(X_{i}\mathrel{\,|\,}\mathbf{A})$ and $\mathbb{P}(X_{j}\mathrel{\,|\,}\mathbf{A})$ with $\hat{f}$ and $\hat{g}$, respectively, by regressing $X_{i}$ on $\mathbf{A}$ and $X_{j}$ on $\mathbf{A}$. A test statistic based on the mean squared prediction error (MSPE) across all observations $x_{i}\in X_{i},x_{j}\in X_{j}$, conditional on all $\mathbf{A}$, is then calculated. The significance level at which to decide dependence must be chosen by the practitioner (see 4.2.1 for further discussion).
No one function or model class is proposed to be ideal for the GCM test. The authors show that the test statistic is asymptotically standard normal if $n$ times the expected product of MSPEs of $\hat{f}$ and $\hat{g}$ approaches 0, the products of the MSPE and regression error of each model converge to $n$, and the expectation of the product of model variances is finite and greater than 0. This result is uniform across all distributions in the null space given additional restrictions on the uniformity of MSPE product convergence to 0 and moment conditions on the expected product and norms of model variances. Any models meeting these criteria for all distributions in the null space can be used in the GCM test to achieve power at any alternative. The authors discuss kernel ridge regression and boosted regression trees as examples; see [52] for proofs, empirical support, and further details. We discuss practical aspects of model selection in light of our causal discovery framework in section 4.
3 Causal discovery for spatiotemporal data
Methods described so far assume all observations are independent and identically distributed, and that there are no confounding variables missing from the system. Data distributed across space and time are often autocorrelated, however, and may depend on unobserved time-lagged and spatial confounders [9, 10].
These obstacles are dealt with differently according to the goal or overarching question asked of causal discovery. Much literature on causal discovery for spatiotemporal data currently focuses on estimating patterns of influence within and/or between a handful of variables on a large spatiotemporal scale [62, 54, 53, 20, 26]. Spatial autocorrelation is often assumed irrelevant due to domain knowledge [54, 53] or removed by estimating spatial factors that aggregate nearby points and are assumed to be effectively independent of each other [62, 20, 30, 48, 57]. While some methods allow for relationships to vary continuously with time [22], others look at temporal snapshots [26]. The estimated causal patterns summarize observed phenomena and can inform extrapolative forecasting models.
Alternatively, a relatively understudied goal of causal discovery for spatiotemporal data is to infer the underlying structures of causal relationships between variables whose observations are distributed and autocorrelated across space and time. This approach imagines that there are latent mechanistic relationships that vary from an underlying central truth across space and/or time. These “latent mechanism,” rather than “latent pattern,” methods are more relevant for fields like ecology, economics, and public health, where interpretability of causal discovery output has policy and management implications (see illustration in Figure 1; [39]).
Figure 1: A simplified illustration of a system for which spatiotemporal latent mechanism causal discovery is relevant. In this system, fires cause instantaneous changes in forest cover across Great Britain, which in turn causes changes in a theoretical species’ abundance at a time lag. Although the variables are spatially distributed, practitioners seeking to understand causes of an observed shift in species abundance will be more interested in the mechanisms of change between variables than the spatiotemporal patterns of each variable.
This paper seeks to address spatiotemporal autocorrelation and confounding relevant to latent mechanism causal discovery, which has yet to receive much attention [6, 29]. In the following sections, we briefly review the literature on time series causal discovery and causal discovery across contexts. We then discuss what causal discovery can learn from methods that address unobserved spatial confounding in causal effect estimation, and how this allows us to extend time series causal discovery to systems with spatial structure.
3.1 Time series causal discovery
Temporal structures necessitate a treatment of autocorrelation and lagged effects. Autocorrelation can arise when values of a variable at one time are caused by values of that variable in the past [15]. This serial dependence can be dealt with by shifting whole time series back by some time lag and including the shifted variables as other points, or nodes, in the system with which to explicitly test for relationships [45]. These shifted time series then allow us to test for relationships between different variables that act at a time lag and would be confounded or at least over-simplified in non-temporal causal discovery [3].
The maximum time lag to consider must be chosen by the researcher as a hyperparameter [47, 36]. The choice is a tradeoff between potentially oversimplifying biological processes or unnecessarily complicating the system and reducing power, as the number of nodes increases from $p$ observed time series to $p+Tp$, where $T$ is the maximum time lag considered.
Despite this increase in the number of nodes and tests, causal discovery for time series can be less complex than methods for i.i.d. data since we assume effects never precede their causes. The computational complexity of time series causal discovery algorithms is polynomial to the number of variables in the worst case rather than exponential (as for the PC algorithm) [47, 46]. Time-series extensions of constraint-based causal discovery are thus focused on how to reduce complexity by best taking advantage of temporal structures [1].
The PCMCI+ algorithm is an example of a popular constraint-based, PC-derived method to efficiently discover time-lagged and instantaneous relationships between time series [44]. It optimizes the PC algorithm for time series by splitting the task of causal discovery into two stages: one for lagged effects, and another for instantaneous. The lagged stage mirrors the PC algorithm, considering only edges between nodes in the “present” ($\mathbf{X}_{t}$) and “past” ($\mathbf{X}_{t-\tau},1\leq\tau\leq T$). Since we assume effects cannot precede their causes, the directions of all edges discovered in the lagged phase are known and flow in the direction of time. It follows that no past node can be a collider, or common cause, of two present nodes as that would imply an edge directed backwards in time. Furthermore, since only colliders can induce dependence, the inclusion of past nodes cannot affect the search for separating sets $\mathbf{A}_{it,jt}\subseteq\mathbf{X}/\{X_{it},X_{jt}\}$ such that $X_{it}\mathrel{\perp\!\!\!\perp}X_{jt}\mathrel{\,|\,}\mathbf{A}_{it,jt}$. When searching for separating sets between present nodes in the instantaneous phase, we can therefore restrict the search space to sets including past nodes adjacent to $X_{i,t}$ and/or $X_{j,t}$ without worrying about potential colliders.
Most assumptions of the PC algorithm are also made by PCMCI+; the time series algorithm assumes causal faithfulness, sufficiency, and stability of relationships across time [44]. Relative to other time series causal discovery algorithms, however, PCMCI+ avoids assumptions on the linearity and distribution of variables [36, 16], and is unique in allowing both time-lagged and instantaneous relationships [47, 24, 1]. Relaxation of the causal sufficiency and stability assumptions have been developed, but at the cost of higher computational complexity and data requirements [21, 8, 49].
PCMCI+ also continues to assume independent noise across the system, albeit now given that temporal autocorrelation and time-lagged relationships have been dealt with explicitly [44]. The algorithm has been shown to be highly sensitive to non-independent noise when data have been collected with spatial structures, even if nodes are included as a pre-specified spatial grid [27]. To use the efficient logic of PCMCI+ for spatiotemporal, latent mechanism causal discovery, we must draw upon methods for spatial causal inference and explicitly extend the algorithm.
3.2 Spatial causal discovery
In latent pattern causal discovery, nodes represent explicit spatial points, and a constraint similar to temporal information can be used to improve efficiency when one assumes causes do not “jump” over locations and must instead propagate through adjacent areas up to a maximum distance threshold [68, 26]. Latent mechanism causal discovery does not have to deal with the same explosion of complexity that comes with making each location its own node, but must grapple instead with how spatial structures may confound or otherwise obscure our ability to learn causal structures.
3.2.1 Joint causal inference
Perhaps more relevant for latent mechanism causal discovery are methods developed for causal discovery across “contexts”, exogenous variables that may impact but are not impacted by system causal relationships. The joint causal inference framework introduced in [25] has been applied to extend PCMCI+ to time series collected across different contexts or compiled from different datasets [13].
Stepping toward spatiotemporal causal discovery, this joint PCMCI+ (J-PCMCI+) algorithm allows for “spatial contexts,” or variables that vary between datasets but are constant over time within datasets. Assuming that contexts are exogenous to the observed system and that no latent contexts confound observed contexts and system variables, J-PCMCI+ can consistently identify the correct system graph. The authors also show through numerical experiments that the use of one-hot encoded spatial “dummy” contexts that correspond to dataset IDs is sufficient to deconfound system variables [13].
The use of these dummy contexts adds as many new variables as locations where data were collected, significantly increasing the complexity of causal discovery. If we understand spatial structure to mean that observations exhibit autocorrelation that decays with increased distance, we should be able to use the (known) distances between observations to more efficiently and precisely deconfound [66]. Here we turn to the literature for causal effect estimation where adjustments for spatial confounding have been relatively better explored.
3.2.2 Latent spatial confounding in causal effect estimation
Spatial confounding of a causal effect comes about when bias is introduced by omission of a variable with spatial structure [10]. Where spatial confounding is known to exist, but cannot be explained by measured variables alone, spatial coordinates can be used as a proxy. As long as the unmeasured confounders can be thought of as measurable functions of space, and the measured variables in the system have non-spatial variation, spatial coordinates may even capture unobservable or unquantifiable confounding variables [51]. Unbiased causal effects can be estimated by controlling for spatial coordinates, given a sufficiently flexible nonparametric model is used. A discussion grounding this treatment of spatial confounders in the instrumental variables literature can be found in [66].
This method is highly sensitive to the spatial scale of unmeasured confounders relative to observed variables [31]. Specifically, the resolution of unmeasured confounders must be coarser than that of system variables; effects are unidentifiable when there are strata of unmeasured confounders for which only one value of a system variable is possible [51]. Given that unmeasured confounders are, obviously, unmeasured, this is an untestable assumption and must be made based on domain knowledge and taken into account in interpretation.
That said, we hypothesize that the adaption of causal effect estimation methods to causal discovery improves the robustness of such methods to assumption violations and weakens conditions for consistency. Causal discovery asks a simpler question (is there a relationship between these variables?) than effect estimation (what is the relationship between these variables?); we expect biases arising from assumption violations and/or small samples more readily and meaningfully alter a point estimate of effect strength than the assessment of relationship existence.
4 Extending time series causal discovery to variables with spatial structure
Spatial confounding violates the assumptions of pre-existing algorithms for causal discovery with time series [45]. Drawing upon the theoretical results of location as a proxy for spatial confounding in causal effect estimation [10], and the logic of exogenous contexts in joint causal inference [13], we propose time series causal discovery can be extended to spatial structures by including location in the approximating models of generalized covariance measure independence tests [52].
4.1 Model choice
As outlined in section 2.2, given some relatively weak conditions on the mean square prediction error, and because the null is restricted when defining the structure of variables’ distributions and approximating models, the GCM test achieves asymptotic power guarantees with good model structure choice [52].
If we imagine in our framework that spatial contexts may cause observed relationships to diverge from an underlying, mechanistically-relevant mean, hierarchical or mixed effects models are a natural choice. Hierarchical generalized additive models (HGAMs) allow these varying relationships to take nonlinear forms and avoid assuming parametric structures [34]. Provided that splines are chosen to maintain smoothness even in high-dimensional settings (true of most default options in popular software like thin-plate splines and with higher-order penalization [64]), HGAMs with sensible choices for response distributions should meet GCM criteria. The use of spatial coordinates as a proxy for unmeasured spatial confounding leads us to specify a distance-based correlation structure, improving model predictive power and reducing computational complexity.
The effect of the amount and location of knots on smoothness as relevant for GCM conditions remains an open question. We intend to explore how misspecification of knots in HGAMs used for the GCM test affects test power through simulation in future work. Initial exploration indicates that the same sensibilities that help us choose knot amounts and locations in predictive modeling are reasonable to ensure appropriate models for the GCM test [40].
4.2 Assumptions
If we accept the utility of the GCM test and are happy enough with the power granted from our model structure choices, we still have to make some causal discovery and spatial assumptions.
Exogeneity of spatial information is necessary, but luckily it is generally supportable in natural systems. This will rely on study design, however; if locations were sampled because they were expected to exhibit certain trends, the spatial exogeneity assumption will not hold. Practitioners should assess the sampling design(s) under which their data were collected, and consider methods to reduce dependence of locations on data values by down-sampling, for example, when necessary. See [61] for a fuller discussion of bias in spatial sampling.
4.2.1 Faithfulness and uncertainty
[…]
  • PC default: Is $X_{k}$ not in the minimum separating set found for $X_{i},X_{j}$?
  • Majority: Is $X_{k}$ not in the majority of relevant separating sets found for $X_{i},X_{j}$?
  • Conservative: Is $X_{k}$ in none of the relevant separating sets found for $X_{i},X_{j}$?
where relevant sets are defined as those containing all variables that have the potential to be common causes of $X_{i}$ and $X_{j}$. Practitioners may choose to be more conservative with regards to faithfulness at the cost of higher complexity.
Most algorithms do not provide any measure of uncertainty in edge assignment and/or direction. Bootstrapping is a natural choice when parametric estimates of variance seem impossible, but requires a large dataset whose structure would not necessarily be disrupted by resampling; this is not the case for spatial time series. A parametric bootstrap may be possible, but either way rerunning an already complex algorithm enough times to generate a reasonable sample would likely be prohibitive.
As an alternative, Petersen et al. (2021) propose reiterating a causal discovery algorithm across many significance levels to capture the “strength of support” for or certainty of each causal relationship. More certain relationships should continue to appear in graphs with stricter thresholds that admit fewer edges [39].
4.2.2 Sufficiency and unmeasured non-spatial confounders
As discussed in [10], only those unmeasured confounders with spatial structure will be accounted for by the inclusion of coordinate proxies. Unmeasured confounders without spatial structure are as yet unaddressed by our proposal.
Among algorithms that allow for latent confounders ( i.e. relax the causal sufficiency assumption) the IC* algorithm is easiest to integrate with our proposal as it only diverges from the PC formula in its edge directing phase [56]. The IC* algorithm returns a marked graph with four types of edges:
  1. 1.
    $X_{i}\mathrel{\xrightarrow{*}}X_{j}$ indicates a genuine causal influence of $X_{i}$ on $X_{j}$
  2. 2.
    $X_{i}\rightarrow X_{j}$ indicates that there is either a genuine causal influence of $X_{i}$ on $X_{j}$ or that there is some latent common cause $X_{i}\leftarrow U\rightarrow X_{j}$
  3. 3.
    $X_{i}\leftrightarrow X_{j}$ indicates a latent common cause $X_{i}\leftarrow U\rightarrow X_{j}$
  4. 4.
[…]
If practitioners are not willing to make assumptions of nonspatial sufficiency, they can swap sufficiency-assuming rules for edge orientation (e.g. PCMCI+ [44]) for IC*’s final phase. Those willing to make such assumptions can enjoy the relatively higher power that comes with causal sufficiency [33].
4.2.3 Stability and temporal stationarity
We make the implicit assumption that relationships between variables do not change over time. J-PCMCI+ allows for temporal contexts that may capture some of that nonstationarity [13], and there is interesting work on breaking up time series into “regimes” under which different relationships between variables may be at play [49, 22]. Work would be needed, however, to incorporate that structure into causal discovery algorithms that are also spatially explicit and to see how large a sample size is needed to achieve power with that much allowed variability.
5 Perspectives and future directions
In this section, we discuss remaining challenges and open questions in the development of a causal discovery algorithm for spatially-structured time series. We then outline specific potential applications and general intended impact of our work.
5.1 Algorithm implementation and evaluation
5.1.1 Computational intensity
One of the largest challenges in developing causal discovery algorithms is computational intensity [19]. Despite choosing models and designing the algorithm to be as time-efficient as possible without losing power, the sheer task of fitting all models to approximate conditional distributions for all pairs of variables conditional on iterative sets of the rest of the system is gargantuan [34].
Parallelization offers a promising path to faster overall algorithms, but is difficult in practice due to the interdependence of stages of causal discovery (i.e., which conditional independence tests are run depends on findings from previous testing). In a nice balance of the efficiencies gained from parallelization and from information sharing across conditional independence tests, Le et al. (2019) proposed parallel-PC, which searches for separating sets for different variable pairs across multiple cores and then synchronizes information across at each iteration [19]. These adjustments significantly reduced runtime compared to the order-independent PC algorithm. Efficiency improved with more cores even on very large datasets.
More complicated algorithms, such as our spatiotemporal proposal, are perhaps less amenable to parallelization. Choices we made in initial algorithm design to improve computational efficiency are often at odds with the notion of dividing distinct tasks across multiple cores. Approximating models for the GCM test can be more efficiently fit by utilizing parallelization features already built into computational packages such as mgcv [65]. Tests of a group of separating sets can be ordered by the test statistic associated with previous tests of similar sets, increasing the likelihood that you find the separating set sooner and need to perform fewer tests overall [47]. The residuals of approximating models can be saved and accessed from temporary storage, eliminating the need to re-fit models for each new GCM test. All of these features require shared information to inform each test and/or access to multiple cores, neither of which is normally possible with parallelization. Further testing will inform whether efficiencies lost in parallelization outweigh efficiencies gained in the spatiotemporal context.
5.1.2 Simulations and sensitivity analysis
Although spatially-structured time series abound in real-world data, their simulation is not so straightforward. This is further complicated by a desire to test the ability of an algorithm to recover potentially nonlinear, non-Gaussian relationships.
Most data for tests of causal discovery algorithms are simulated from the corresponding structural causal models of the “true” graph (see definition in section 2). Aspects of that simulation process create imbalances in variation between variables that causal discovery algorithms can “game” to achieve unrealistic performance [42]. The additive noise at each level of the SCM compounds down the causal order such that ancestral variables have smaller marginal variances than their descendants. Causal discovery algorithms can then take advantage of the predictable variability structure to identify graphs more successfully than they could for standardized or randomly varying systems. This issue is more relevant for methods that exploit asymmetries in noise than constraint based methods, but standardization and/or careful model design are needed to avoid inflating discovery performance on all simulated data sets.
Agent-based models (ABMs, also known as individual-based models in some fields) may provide a more appropriate method of simulating data to benchmark and check sensitivity of causal discovery algorithms. In ABMs, complex system properties arise from individual agents’ interactions with their environment, a philosophy that nicely mirrors the goal of latent mechanism causal discovery [12]. Guides for ABM design suggest visualizing the models with influence diagrams [12] and an emerging literature calls for use of causal discovery in ABM validation [17, 67]. Because noise arises from the randomness of agents’ behavior rather than by sequential addition, data simulated by ABMs are not subject to the marginal variance issues described for structural causal models.
All causal discovery algorithms rely on untestable assumptions. Practical users will be interested in how sensitive any proposed algorithm is to violations of those assumptions. In our simulations, we will want to assess sensitivity of the algorithm to biased location sampling, unmeasured nonspatial confounding, and temporal nonstationarity (section 4.2). Once again, ABMs provide a convenient simulation framework. One could alter, for example, the spatial scale of the environment or the temporal scale at which agents act to check for algorithm robustness to scale mismatch. More work is needed to evaluate if the potential of ABMs for robust causal discovery benchmarking is worth the consideration and computational power needed to simulate them [12].
5.1.3 Causal discovery performance metrics
The literature on evaluating the performance of causal discovery algorithms is sparse relative to that for building them. CauseMe, an online database for sharing and comparing causal discovery algorithms, uses machine learning classifier metrics such as true positive rate, false positive rate, F1, and area under a receiving operator curve to compare performance of a wide variety of algorithms on benchmark datasets [45]. While reasonable at face value, Petersen (2024) has shown that the expected F1 score of a random graph increases monotonically with the number of estimated edges; comparing graphs with different numbers of edges via F1 score (or other scores based on recall and precision) will bias towards more connected graphs [37]. Instead, Petersen suggests comparing graphs produced by an algorithm to a distribution of graphs created via random assignment of the same number of edges. She proposes an associated hypothesis test for differences between estimated graphs and random edge placement.
Other approaches describe metrics for testing the distance between graphs produced by an algorithm and the ground truth. Structural intervention distance describes the similarity of inferences one would make from estimated vs. true graphs [35]. It counts the number of changes one would have to make to an estimated graph to ensure that the structural causal model built from it matched that of the true graph. A recent extension to structural intervention distance allows for evaluation of marked, partial graphs like those output by IC* [14].
None of these methods are immediately applicable to the kinds of graphs produced by spatiotemporal causal discovery. The discussed options weigh all differences from the “true” graph evenly, but when considering the interpretation of time series causal discovery, some relationships are arguably “more wrong” than others. For example, in the system illustrated in figure 1, it would be “more wrong” relative to interpretation if an estimated graph had no connection between forest cover and species abundance than if it had estimated that as an instantaneous relationship. Qualitative comparisons that ask how and where interpretation or conclusion-making has gone wrong may be more practically useful. These questions can be asked in comparisons with both true graphs (for simulations) and/or a distribution of graphs produced by random edge placement [37], and seem for now the best option to evaluate performance of spatiotemporal causal discovery algorithms.
5.2 Applications and impact
Studies that sought to use current methods to identify underlying mechanisms of cause and effect from observed time series with spatial structure have come up short. Those who employed i.i.d. time series algorithms such as PCMCI+ to analyze spatially structured data reported poor performance and lack of interpretability [27, 23]. Other studies used spatial convergent cross mapping, a method that discovers pair-wise, quasi-causal links between independently replicated time series [43, 50, 7, 2]. These studies produced evidence of potential causal relationships between pairs of variables, but were not able to combine these relationships to holistically understand the larger system. Neither approach addressed spatial confounding or used the known spatial structure of their data. Our proposal, though not yet stress-tested or benchmarked, takes an important first step towards achieving causal discovery of mechanisms from spatiotemporal data.
We expect work in this area to have impact outside of statistical theory. The visualization of cause-and-effect relationships in a graph (e.g. figure 1) allows non-specialists to engage in analysis, interpretation, and decision making. In ecology and environmental sciences, causal graphs interface nicely with existing visual policymaking methods. The United States Environmental Protection Agency advocate for the use of qualitative influence diagrams (QIDs) in describing complex systems for environmental policymaking [4]. QIDs visually map controlling factors and impacts to help policymakers decide on interventions and evaluate downstream effects. Methods exist for deriving conditional distribution information from QIDs, and empirical causal graphs can easily be read as influence diagrams [63]. Further examples of graphical methods in sustainable development can be seen in [60]. We intend that future application of spatiotemporal causal discovery be accompanied by nontechnical reports that leverage the legibility of causal graphs.